OpenStudy (anonymous):

What is the solution set of |2x – 3| – 6 = –1? A. 5 or -5 B. 4 or 1 C. 4 or -1 D. No solution

OpenStudy (anas.p):

\[\left| 2x-3 \right|-6=-1\]\[\left| 2x-3 \right|=-1+6\]\[\left| 2x-3 \right|=5\]Square both sides.\[(2x-3)^{2}=25\]\[4x^{2}-12x+9=25\]\[4x^{2}-12x+9-25=0\]\[4x^{2}-12x-16=0\]THEN SOLVE QUADRADICALY. You get the answer as 4,-1 So c is your answer